OpenStudy (anonymous):

The second side of a triangle is 8 inches less than the first side. The third side is 14 inches more than the first side. The perimeter of the triangle is 63 inches. find each side of the triangle. help please...

OpenStudy (anonymous):

x+(14+x)+(x-8)=36 try using this its gonna help

OpenStudy (anonymous):

can you please explain on how you got that equation?!

OpenStudy (anonymous):

ok so i used x as the unknown for the question and so one is going to be x and the side with 8 inches less than the first side will be x-8 and the side with 14 inches more than the first side will be 14 +x....does it make sense?

OpenStudy (anonymous):

ohhhh! okkk

OpenStudy (anonymous):

i just don't really get how to simplify it =/

OpenStudy (anonymous):

like its x+(14+x)+(x-8)=63

OpenStudy (anonymous):

oh do u know how to solve for x ?

OpenStudy (anonymous):

and you multiply x and 14 and you keep going..

OpenStudy (anonymous):

sorry my computer froze when i was answering ure question

OpenStudy (anonymous):

no problem thank you for helping my wiiht this question in the first place!

OpenStudy (anonymous):

like can you teach me step by step on how to solve it please? =/

OpenStudy (anonymous):

yeah sure ..so add the like terms first thats the x's and then subtract 14-8

OpenStudy (anonymous):

thats going to be 3x +6=36 then u take six to the other side and ure going to be left with the equation 3x=36-6 right?

OpenStudy (anonymous):

so x equals 19?

OpenStudy (anonymous):

x equals 10

OpenStudy (anonymous):

but the perimeter is 63 not 36..

OpenStudy (anonymous):

oh sorry yeah i c=got confused back yeah its 19

OpenStudy (anonymous):

so then substitute the 19 with x in the equation

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

wait umm can you help please help me answer one more question?!

OpenStudy (anonymous):

oh sure just post it

OpenStudy (anonymous):

okk! thank you so muchh!

OpenStudy (anonymous):

do i post it here or on another box?

OpenStudy (anonymous):

yeah just close this question then u'll get room for posting the new one..

OpenStudy (anonymous):

ok thankks